OpenStudy (anonymous):

MEDAAL AND FAN What is the surface area of the triangular prism shown? 860 ft 2 920 ft 2 1,040 ft 2 760 ft 2

OpenStudy (anonymous):



OpenStudy (anonymous):

@Mr.ClayLordMath

OpenStudy (anonymous):

@ajprincess @hartnn @dpasingh

OpenStudy (anonymous):

hmm i haven't solved that for while i have to review this

OpenStudy (anonymous):

hartnn can you help her?

hartnn (hartnn):

if you count surfaces, there are 2 right triangles, 2 rectangles and one base can you find area of any of these ?

hartnn (hartnn):

the base is also a rectangle , if you could observe...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

area of a triangle=1/2 *bh

OpenStudy (anonymous):

so i need to do what?

OpenStudy (anonymous):

what formula? how can I solve it?

hartnn (hartnn):

let me take if further, 2 right triangles, height =15, base = 8 Area of a triangle = 1/2 * base * height All 3 Rectangles have different dimensions Area of rect = width * length

OpenStudy (anonymous):

good...so i need to multiply the base heigh lenght and width

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so your total area=2*area of triangle+area of r1 +aarea of r2 + area of r3. where r1,r2 and r3 are the three rectangles

hartnn (hartnn):

first find area of 2 triangles = 2* (1/2 * 15 * 8)=...? then find dimensions of each rectangle, one rect has , w=15,l=20, Area = 15*20 =.. ? other has, w=17, l=20, Area = 17*20 = ... ? lastly, smallest one, the base, has w=8,l=20, Area = 8*20 =...?

hartnn (hartnn):

Add all these to get total surface area

OpenStudy (anonymous):

ok just a sec

OpenStudy (anonymous):

1= 120 dimensions=340?? area=300 area2=340 area3=160

OpenStudy (anonymous):

correct?

hartnn (hartnn):

all are correct :) just add them

hartnn (hartnn):

120+160+340+300

OpenStudy (anonymous):

ok...thanks answer = 920?

OpenStudy (anonymous):

and i'm already your fan

hartnn (hartnn):

correct! for both :)